jennifer is 6 years older than sue. in 4 years, she will be twice as old as sue was 5 years ago. find their ages now.
Answer by oberobic(2304)

J = Jennifer's age
S = Sue's age
J = S+6
.
In 4 years...
J+4
.
J will be twice as old as S was 5 years ago
J+4 = 2(S-5)
J = 2S - 10 - 4
J = 2S -14
.
Substitute
.
S+6 = 2S-14
-S = -20
S = 20
J = 20+6
J = 26
.
Check.
In 4 years:
J+4 = 26 + 4 = 30
...
Jennifer will be twice as old as Sue was 5 years ago.
5 years ago Sue was 20-5 = 15 years old
2 * 15 = 30
...
Correct.
.
Answer:
Jennifer is 26, and Sue is 20.
.
Done.